Mathematics 242 – Principles of Analysis Definitions Quiz 2 – February 7, 2014

Your Name:

A) (5) What does the Principle of Mathematical Induction say? (Just give the statement, not the proof.)

B) (5) Express the set {|x + 3| | x ∈ (−2, 4)} as a union of one or more open intervals.